A wire of length 3.45 m and mass 13.70 g is under a tension of 1.25 N. A standing wave has formed which has seven nodes excluding the endpoints. Calculates the frequency of this wave?




1
Expert's answer
2021-11-29T11:40:14-0500

The number of harminic is "n = 7+1 = 8", since the first harmonic has 0 nodes excluding the endpoints. The frequency is given as follows:


"f_n = \dfrac{nv}{2L}"

where "L = 3.45m" and "v" is the speed of the wave. "v" is given as follows:


"v = \sqrt{\dfrac{FL}{m}}"

where "F = 1.25N, m = 13.70g = 0.01370kg". Thus, obtain:


"f_n = \dfrac{n}{2L}\sqrt{\dfrac{FL}{m}} =\dfrac{n}{2}\sqrt{\dfrac{F}{Lm}} \\
f_8 = \dfrac{8}{2\cdot}\cdot \sqrt{\dfrac{1.25}{ 3.45\cdot 0.01370}} \approx 20.6Hz"

Answer. 20.6 Hz.